### v1

# Deformable Cluster Manipulation via Whole-Arm Policy Learning

Abstract
Manipulating clusters of deformable objects presents a substantial challenge with widespread applicability, but requires contact-rich whole-arm interactions. A potential solution must address the limited capacity for realistic model synthesis, high uncertainty in perception, and the lack of efficient spatial abstractions, among others. We propose a novel framework for learning model-free policies integrating two modalities: 3D point clouds and proprioceptive touch indicators, emphasising manipulation with full body contact awareness, going beyond traditional end-effector modes. Our reinforcement learning framework leverages a distributional state representation, aided by kernel mean embeddings, to achieve improved training efficiency and real-time inference. Furthermore, we propose a novel context-agnostic occlusion heuristic to clear deformables from a target region for exposure tasks. We deploy the framework in a power line clearance scenario and observe that the agent generates creative strategies leveraging multiple arm links for de-occlusion. Finally, we perform zero-shot sim-to-real policy transfer, allowing the arm to clear real branches with unknown occlusion patterns, unseen topology, and uncertain dynamics.
Videos are available @ https://youtu.be/hG-zYoesX5k
I INTRODUCTION
Learning to manipulate deformable objects poses a formidable challenge in robotics [1][2]; however, operating with clusters of deformables, such as cable bundles, multi-branch tree canopies, and cloth piles, is significantly harder. Although imprecise physics knowledge [3] and sensor uncertainties are contributing factors, the principal challenge lies in inferring a compact, low-dimensional state abstraction and the temporal state evolution, based on noisy perception from a high-dimensional ground truth [4]. Furthermore, with deformable clusters, tiny differences in contact locations, material properties, and initial configuration among the individual cluster constituents (e.g. between thicker and thinner plant stems) give rise to non-linear resistance forces and divergent motion trajectories upon interaction. The resulting chaotic dynamics render model-based learning infeasible, while model-free solutions have high sample complexity, and therefore, the compactness and generation efficiency of the state representation gather significance [5].
Figure 1: Overview: (1) Segmented point clouds corresponding to the clustered deformable branches, clearance region, and the robot are captured from the scene. (2) Distribution embeddings representing global scene are generated via kernel mean operator. Additional features include neighbourhood point features, robot sensor metrics, and proprioceptive contact indicators. (3) RL training with domain randomised geometry and dynamics on Isaac Gym parallel simulator. (4) Inference & zero-shot transfer to the real world aided by Grounding DINO, SAM-HQ, and Cutie frameworks.
Novel deformable works manipulating cables [6], dough [7], and clothes [8] are all planning approaches; furthermore, they exclusively focus on end-effector-based control. In a cluster context, such grasp-based solutions tailored to individual deformable members become prohibitively long. On the contrary, a multi-link manipulation strategy substantially escalates the contact complexity, necessitating additional perception modalities apart from vision. However, typical auxiliary sensors for whole-arm contact detection are either expensive [9] or too slow [10] for low-latency applications.
Addressing these challenges, we propose a model-free, non-prehensile, multi-modal, reinforcement learning approach (Fig. 1) operating on the cluster as a whole with full-body contact, but without external tactile sensing. We use a distributional representation [4] of the cluster state generated by embedding point cloud samples into a Reproducing Kernel Hilbert Space (RKHS), without key-point intermediaries. Compared to traditional point cloud feature extractors, a distribution embedding is notably lightweight, enabling faster training and real-time reinforcement learning (RL) feedback. It eliminates the need for points to track identical cluster locations, is invariant to permutations, and can handle variable input sizes resulting from cluster members moving out of camera view between time steps. Further, the observations include a whole-arm contact classifier [11] output from proprioceptive measurements to flag contact. Finally, we address the high sample requirements of on-policy RL by leveraging the parallelism of physics simulators for policy training and transferring it zero-shot to a real-world setting with completely unseen geometry and dynamics.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d)
Figure 2: (a) Manual pruning of overhanging branches impeding power lines, from [12] (b) Network cables occluding autonomous server inspection (c) Our laboratory setting mimicking the power line clearance scenario (d) Simulation environment in Isaac Gym with a parametric L-system ternary structure.
This work focuses on a specific subclass of problems, aiming to de-occlude local regions of the deformable cluster; for instance, to clear a cylinder-shaped region among tree branches, by manipulating a few of them together. Illustrative applications of such a skill abstraction include:
  1. 1.
    Clearing Power Lines: Hazardous tree branches encroaching on overhead power infrastructure is a leading cause of forest fires and outages [13][14]. Pruning foliage near overhead lines (Fig. 2a). is dangerous for humans due to risks like falls, shocks, and tool accidents [12]; therefore, robotic assistive systems that can clear the branches away from the lines before pruning can enhance safety while reducing outages. In this instance, the cylindrical clearance region can be thought of as concentric to the physical power line.
  2. 2.
    Autonomous Inspection: Modern electrical and networking infrastructure, such as a data centre, is cluttered with deformable electrical and network cables obstructing inspection cameras (Fig. 2b). Autonomous monitoring robots must flexibly manipulate cable bunches to restore camera line-of-sight for back panel inspection and anomaly detection. Here, the clearance region may be a virtual cylinder with the inspection camera on one face and the target panel on the other.
  3. 3.
    Agricultural Exposure: Modern harvesting systems suffer from significantly lower success rates with partially (50%-75%) and fully occluded (5%) fruits [15]. Manipulating plant stems for exposing fruits or a diseased region to an external sensor requires a line-of-sight restoration through multi-branch clearance.
Among these, the first problem presents the most significant challenges due to the additional constraints imposed by the rigid power line on the robot’s motion, and thus serves as the primary focus of this paper in both simulation and real settings, illustrated in Fig. 2 (c) & (d). While our aim is not to provide an industry-grade system, it is noteworthy that no fully autonomous solutions currently exist for the power line problem. Additionally, we provide simulation videos for the agricultural exposure task to emphasise the generality of our approach. Formally, our contributions are as follows:
  1. 1.
    A multi-modal, whole-body contact policy framework to manipulate deformable clusters, learned from both point clouds and proprioceptive touch detection inputs.
  2. 2.
    An efficient distributional state representation of the complex deformable cluster geometry, with low computational overheads, in an RL policy learning context.
  3. 3.
    A novel context-agnostic occlusion heuristic and reward strategy generalizable to various exposure applications to clear deformables from a target region.
  4. 4.
    A specific implementation to clear power lines of overhanging tree branches and zero-shot sim-to-real transfer that handles unseen branch topology and displays novel clearance patterns in simulation and real.
Subsequent sections specifically focus on the power line clearance problem; however, the methodology is readily adaptable to others. We encourage readers to view the extensive supplementary videos to see our approach in action.
II RELATED WORK
For deformable manipulation with high-dimensional object geometries, point clouds offer a rich sensor-agnostic scene mapping. Notable 1D implementations for tracking [16] and learning dense depth object descriptor features [17] leverage point representations. Point clouds have been used to learn policies with 2D and 3D deformables as well for cloth manipulation [8], liquid scooping [18], and dough cutting [7]. On the policy front, these deformable works use either Motion Planning with primitives [8], Task and Motion Planning [7], or Behavioural Cloning [18][17]. In contrast, model-free reinforcement learning works on deformables like ours are scarce, largely due to poor sample efficiency. With rigid body point clouds, model-free policies have been proposed [19][20] for interactive tasks, including sim-to-real transfer. Relevant to our setup, a key study with rigid bodies [21] found that RL with 3D point clouds outperforms its 2D image counterpart when object-object relationship encoding is crucial for the task.
Post-scene capture, common deformable representation strategies include keypoints [16], graphs [22], and dense-descriptors [17]. Particularly relevant to our work, the deformable state can be represented as as a distributional embedding with a kernel mean operator [4]. They tackle a real-to-sim inference problem with an RKHS-net layer using a small number of keypoints ($\approx 10$) extracted from images. In contrast, we project entire point clouds ($\approx 512$) of multiple objects to RKHS and use the embeddings directly as RL observations in a sim-to-real context.
Our policy enables the arm to contort itself into shapes, such as hooks and props, leveraging the whole arm to push away the occlusion. While such formations are less prevalent, contact-aware reaching has been achieved with external full-body tactile sensors, for rigid clutter [9], or with internal sensors for tree branches [11]. However, we go beyond simple reach to sophisticated clearance tasks and implement a multi-modal policy integrating touch and vision feedback.
Among deformables, interacting with plant structures is substantially harder owing to the non-uniformity in geometry, dynamics, and visual features. [23] proposes a method to gently grasp and move a leaf aside to estimate the pose of hidden fruits. On the other hand, [24] pushes away multiple leaves simultaneously with the end-effector to reveal the space behind them using large amounts of real-world data. Both works assume soft leafy resistance, which is typically addressed in farms with simpler tricks, such as blowing compressed air [25] to separate leaves and fruits. From a learning perspective, [26] and [27] employ vision-based RL for reaching vine pruning locations and for exposing fruits, respectively; however, both assume known plant topology while focusing on the end-effector motion. In comparison, our approach is designed to accommodate stiff resistance from multiple branches, uses the whole arm to maximise applied torques, and is trained in simulation without any digital twins, plant mesh similarity, or real-world data. Finally, none of the aforementioned works are designed for the hybrid rigid-deformable problem of power line clearance.
III APPROACH
III-A Deformable Simulation
Realistic simulation of the intricate tree-branch geometry is challenging; fortunately, this is well-studied in botanical literature under the L-system paradigm [28], predominantly for visualisation. An L-system borrows from formal grammar theory to model plant morphology by applying a recursive rule collection, starting from an axiom, mimicking the branching patterns. Recently, [11] has shown that this paradigm can be exported to 3D physics simulators, replete with realistic dynamics, by representing branch segments as rigid cylindrical links connected by revolute joints, amenable to robotic interaction. In this setup, branch deformations are modelled with a mass-spring-damper [29] actuated with proportional-derivative controllers. We borrow the recursion rules, growth attributes, and L-system parameters from [11]; however, we extend it with spherical joints to enable multi-axis deformation and enhance compliance for clearance, refer Fig. 2(d). We run thousands of such tree models in parallel on the distributed physics simulator Isaac Gym [30].
III-B Policy Learning
We formulate the de-occlusion task as a discrete-time Markov Decision Process (MDP), where an agent learns a stochastic policy $\pi_{w}(a|o)$; parameterised by weights $w$. An MDP is defined by $\langle S,A,P_{a},R_{a},\beta\rangle$, where $S$ is states, $A$ is actions, $P_{a}$ and $R_{a}$ are the state transition model and the reward from action $a\in A$, and $\beta$ is a discount factor. The agent aims to maximise $\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\beta^{t}R_{a}(s_{t},a_{t})\right]$, the expected discounted rewards over $T$ steps. Crucially, the agent only has access to noisy observations $o\in O$; for instance, point clouds from the camera or robot sensor measurements rather than the true state. Our algorithm choice is Proximal Policy Optimisation [31] from RL Games [32].
III-C Kernel Mean Embedding
In this section, we briefly examine distribution embedding and its spectral approximation before applying it to deformables; for a thorough explanation, see [33] and [34].
An RKHS is a Hilbert space of functions uniquely determined by a positive definite kernel function $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, that guarantees both an implicit feature mapping $\varphi:\mathcal{X}\to\mathcal{H}$ from the input space and the higher-dimensional target space $\mathcal{H}$. Akin to point mappings, a kernel mean operator lifts a probability distribution $\mathbb{P}$ to a single mean function in $\mathcal{H}$, however, without any loss of information for a characteristic kernel such as RBF: $k(x,x^{\prime})=\exp(-\frac{||x-x^{\prime}||^{2}}{2\gamma{}^{2}})$. We imagine $x$ to be points from the 3D space of point clouds. The lifted kernel mean is defined as:
[…]
Intuitively, formulation (1) could be viewed as the average kernel similarity of parameter $x$ to all domain points according to the distribution $\mathbb{P}$ in RKHS. When the underlying distribution is not explicitly known, which often is the case, an empirical kernel mean $\overline{\mu_{\mathbb{P}}}$ can approximate the true $\mu_{\mathbb{P}}$ from just the i.i.d samples $\{x_{1},\dots,x_{N}\}$ drawn from $\mathbb{P}$.
$\overline{\mu_{\mathbb{P}}}:=\frac{1}{N}\sum_{i=1}^{N}\varphi(x_{i})=\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\langle\varphi(x_{i}),\varphi(x_{j})\rangle_{\mathcal{H}}$ (2)
While the inner product $\langle.,.\rangle_{\mathcal{H}}$ in eqn. (2) can indeed be computed with a finite-dimensional gram matrix, leveraging the reproducing property $\langle\varphi(x_{i}),\varphi(x_{j})\rangle_{\mathcal{H}}=k(x_{i},x_{j})$, but this is still problematic for large N. On the other hand, representing kernel functions through their spectral representations, particularly with Random Fourier Features (RFF) [34], significantly reduces the computational overhead. Through Bochner’s theorem, any kernel can be defined as a Fourier transform of a non-negative measure, subject to a stationarity constraint $k(x,x^{\prime})=k(x-x^{\prime})$.
A Monte Carlo estimate of the kernel can then be written as $k(x,x^{\prime})\approx\frac{1}{R}\sum_{r=1}^{R}\cos({\omega}_{r}^{\top}(x-x^{\prime}))$ where the frequencies $\omega_{r}$ are sampled from a spectral density, $\omega_{r}\sim p(\omega)$. The above estimate becomes computationally efficient for $R<<N$, which we exploit. If $k(x,x^{\prime})\approx\hat{\varphi}^{\top}(x)\hat{\varphi}(x^{\prime})$, the empirical kernel embedding can be further simplified as:
$\overline{\mu_{\mathbb{P}}}:=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\sqrt{R}}\begin{bmatrix}\cos(\omega_{1}^{\top}x_{i}),\sin(\omega_{1}^{\top}x_{i}),\dots,\sin(\omega_{R}^{\top}x_{i})\end{bmatrix}.$
We choose $R=16$ and use the RBF kernel, which results from the Fourier transform of a Gaussian $p(\omega)\sim\mathcal{N}(0,1)$.
III-D Distribution Embedding for Deformables
Given a segmented point cloud $P_{t}\in\mathbb{R}^{N\times 3}$ at time-step $t$, our spatial abstraction assumes it as samples from an underlying distribution, regardless of the deformability. The points in 3D coordinate space are elevated to the feature space $\mathcal{H}$ with an RFF approximated kernel mean, as described in Section III-C, to form the key constituent of the observation space $O$. We capture four independent point clouds of varying sizes: $P^{rob}$, $P^{clr}$, $P^{wbr}$, $P^{zbr}$, representing the robot, clearance region (i.e. a slice of the power line), the whole tree, and a zoomed-in version of the branches near the occlusion. The robot points are sampled from the URDF mesh based on link positions at each time step. Points representing the cylindrical clearance region are sampled from the cylinder surface, given its fixed position and orientation. In contrast, camera sensors provide the deformable and time-varying points of the occluding branches.
There are several advantages to such a representation. Firstly, although the explicitly mapped feature locations are unknown, the agent can exploit the relative similarity of embedded points in RKHS, for instance, to minimise the foliage-power line entanglement or maximise the gripper-branch proximity, resembling the ‘kernel trick’ common in machine learning literature. Secondly, a distributional interpretation is invariant to point permutations and robust to the high noise content characteristic of streaming point clouds. Third, unlike conventional point cloud feature extractors like PointNet++ [35], kernel embeddings are resilient to variations in the size and structure of input point clouds, amplified by the arm, branches and the power line frequently occluding each other. Finally, and most crucially, this approach is significantly faster, reducing training time and enabling real-time inference, as our experiments demonstrate. This speed advantage stems from formulation in Section III-C, where the feature map $\hat{\varphi}(x)$ can be derived with a single matrix product of the input points $x$ and the pre-computed Fourier coefficients ${\omega}_{r}^{\top}$ of complexity $\mathcal{O}(N\times R)$, followed by a sin/cos transformation, and an additional row-average to obtain the embedding $\overline{\mu_{\mathbb{P}}}$, a highly efficient pipeline scalable to thousands of environments, amenable to GPU parallelisation.
III-E Occlusion Heuristic
A key consideration in the de-occlusion problem is determining an effective yet simple measure to capture the temporal progression in occlusion level from manipulation. Common metrics include visible surface ratio, pixel counts [36] (e.g., red fruit pixels in a green foliage background), or labelled occlusion levels. However, these metrics depend either on prior knowledge of the target geometry or partial observability at start, and they may not be smooth enough for an RL agent. In contrast, our occlusion heuristic $h_{t}\in[0,1]$ is independent of target characteristics or even its presence.
Given segmented point clouds $P^{(1)}\in\mathbb{R}^{N_{1}\times 3}$, $P^{(2)}\in\mathbb{R}^{N_{2}\times 3}$ of two entangled entities, the obstructing branches $P^{zbr}$ and a virtual clearance region $P^{clr}$ in our case, the $h_{t}$ is computed as follows. First, we define a pairwise distance matrix $D\in\mathbb{R}^{N_{1}\times N_{2}}$:
$D_{i,j}=\|p^{(1)}_{i}-p^{(2)}_{j}\|_{2}\text{ for }i=1,\dots,N_{1};\ j=1,\dots,N_{2}.$
Let $\mathcal{S}\subset\{1,\dots,N_{1}\}\times\{1,\dots,N_{2}\}$ be the indices set corresponding to the $k$ smallest distances in $D$, i.e., the $(i,j)$ indices of the $k$ nearest-neighbor point pairs, and $d_{th}$ be a distance threshold, then $\mathcal{D}_{k}:=\left\{D_{i,j}\mid(i,j)\in\mathcal{S}\right\}$ and
$h(P^{(1)},P^{(2)}):=\frac{1}{k}\sum_{d\in\mathcal{D}_{k}}\mathbb{I}(d<d_{\text{th}}).$
Intuitively, it is the fraction of point pairs (from different groups) that breach a proximity threshold to the total number of pairs in the neighbourhood, where both the threshold $d_{th}$ and the neighbourhood size $k$ are tunable. In general, the larger the number of branch segments covering the clearance region, the higher the $h_{t}$. We set $k=200$ and the threshold to be the clearance region cylinder diameter $d_{th}=10$cm. We emphasise that the clearance region has no shape constraints due to its point representation; for instance, it can be defined to de-occlude an expected curved end-effector trajectory of a second manipulator in a dual-arm setting.
III-F Observation & Reward Space
TABLE I: Observation Feature Groups
Group Feature Shape
Proprioceptive joint pose: ${q}$ 6
joint velocity: $\dot{{q}}$ 6
ee quaternion: ${r}_{\text{EE}}$ 4
KME whole branches: $\overline{\mu_{\mathbb{P}}}[P^{wbr}]$ 16
zoomed-in branches: $\overline{\mu_{\mathbb{P}}}[P^{zbr}]$ 16
clearance region: $\overline{\mu_{\mathbb{P}}}[P^{clr}]$ 16
robot: $\overline{\mu_{\mathbb{P}}}[P^{rob}]$ 16
Touch contact indicator y/n: $\mathbb{I}(\|F_{t}\|_{2}>f_{u})$ 1
Local PC ee-branch dist: $\mathcal{D}_{k=5}(P^{ee},P^{wbr})$ 5
safety breach indicator y/n 1
Occ Heuristic $h_{t}(P^{wbr},P^{clr})$ 1
All Features 88
Our observation space $O$ consists of five feature sets, listed in Table I. First, we construct two distinct feature sets from point cloud information: the KME features described in Section III-D, which characterise global structure, and a local feature group, which focuses on nearby point distances. The latter includes the Euclidean distance of the end-effector to the closest $k=5$ branch points and a safety breach indicator to reduce the robot-line contacts, computed as:
$\delta_{sm}=\mathbb{I}\left\{\frac{1}{|\mathcal{D}_{5}|}\sum_{d\in\mathcal{D}_{5}}d<d_{\text{sm}}\right\};\ \mathcal{D}_{5}=\mathcal{D}_{k=5}(P^{rob},P^{clr}).$
As the branches overhanging the power line already act as conductors [14], the robot touching the power line in itself does not compromise the validity of our approach; nevertheless, a safety margin $d_{sm}=4$cm is introduced to reduce the wear and tear to the line through reward shaping.
A novel aspect of our method is its multi-modality, attributable to a whole-arm touch detector. Recently, [11] has shown that in the absence of expensive tactile sensors, time-series sliding window torque measurements: $[\tau_{t},\tau_{t-1},\tau_{t-2},...,\tau_{t-(m-1)}]$ over the last $m$ time steps along with other proprioceptive features, can be used to build a classifier trained to detect contact ‘bumps’ at the current time-step $t$, from labeled touch data. While [11] is trained with touch features for a simple reach task, we extend that approach and demonstrate that the touch classifier is useful even in the presence of vision and can aid complex manipulation tasks. During simulation, a proxy classifier $\mathbb{I}(\|{F_{t}}\|_{2}>f_{u})$ computes the force threshold $f_{u}$ breaches by the robot links $l$, leveraging the net contact force $F_{t}\in\mathbb{R}^{l\times 3}$ exposed by Isaac Gym. The full body contact awareness from this touch detector is crucial for the arm to leverage its multiple links (e.g. forearm and wrist) for manipulation, to supplement the partial, noisy, and time-varying point clouds.
Our simple reward structure listed below has three components: a) a clearance reward $r_{h}$ to encourage manipulator towards lower occlusion scores $h_{t}$, b) a smoothness reward $r_{q}$ on the joint velocities $\dot{q}$, and c) a bonus for the time-steps when the arm links are clear of the power line safety margin,
$r_{h}=\left[\frac{1}{1+h_{t}^{2}}\right]^{2},\ r_{q}=-\sum_{j=1}^{6}\frac{\dot{q}_{tj}^{2}}{100},\ r_{sm}=0.4\cdot\mathbb{I}_{\delta_{sm}=0}.$
III-G Domain Randomisation
This work aims to generalise a learned manipulation policy to deformable scenarios with unseen geometry, undetermined dynamics, and devoid of chromatic affinity, by not relying on digital twins [27], parameter inference [29][4], and RGB images respectively during the training phase. Instead, we perturb simulation parameters, for example, L-system structural traits such as branch divergence angle and segment elongation rate, and dynamics parameters such as spring stiffness, damping and friction, among others. In particular, we sub-sample and permute all training point clouds while adding Gaussian noise during each time-step.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d) (e)
Figure 3: Terminal state poses: (a)(b) Simulation trajectory states showing the whole arm being utilised to shield the power line. (c)(d) & (e) Similar real strategies using various arm links for clearance.
IV EXPERIMENTAL SETUP
IV-A Simulation Setting
For the power line clearance scenario illustrated in Fig. 2(d), we run 6144 Gym environments in parallel, each of which contains a tree with deformable branches (in green), a rigid power line (in red), and a Kinova arm actuated via velocity control of its six link joints, omitting the end-effector. Each tree is formed by drawing from a Gaussian distribution ($\sigma=0.1$) representing the L-system morphology parameters to ensure high diversity in topology but sufficient physical feasibility, in addition to arbitrary trunk rotation to expose the agent to varying branch occlusion patterns. Furthermore, the power line location is randomised, but constrained to the reach region of the gripper and semi-parallel (within ±$\pm 15^{\circ}$) to the axis connecting the trunk and the robot base. Finally, we perform all experiments with an average $h_{t}\in[0,1]$ of at least $0.7$ across all environments, noting that occlusion score can increase as well from the arm pushing more branches closer to the power line.
A critical challenge within this framework is the contact explosion resulting from the combinatorial interactions among the multi-link arm, multi-branch tree, and the power line, exacerbated by a large number of training environments. This issue, common in contact-rich parallel simulations, can quickly overwhelm the simulation constraint solvers, triggering object inter-penetrations and CUDA memory overflows. It has been extensively studied by [37] and [38], which suggest complex CUDA-level penetration checkers and mesh-level contact binning, respectively. In contrast, we take a shortcut to contact reduction at the cost of a marginal drop in test performance. Specifically, during training, we allow the branches (but not the robot) to penetrate the power line through collision masking, thereby removing one key set of interactions. On the other hand, during simulation tests, all contacts are in place to accurately reflect real-world, but the environment count is set to a low value, 512 per test dataset. The technique mentioned above to allow for slight differences in training and test environments, along with careful tuning of the Temporal Gauss-Seidel (TGS) contact solver parameters, enables us to run large-scale training required for RL without compromising representational accuracy.
TABLE II: Feature Ablation
Description $d_{\text{max}}$ Train Rew Trials Test Rew Test SROcc DropTouchSteps in Succ
(a) Proprioceptive + Heuristic $h_{t}$ 1495.6 1536 1413.6 ± 28.7 36.5 ± 2.3 18.0 ± 1.3 2.8 ± 0.3 237.7 ± 19.2
(b) (a) + Local PC Features 0.02 1531.7 1536 1486.5 ± 49.1 43.8 ± 4.5 25.2 ± 3.6 2.1 ± 0.4 258.3 ± 32.1
(c) (b) + Touch Feature 0.02 1606.6 1536 1506.1 ± 28.5 46.1 ± 2.6 27.9 ± 4.8 2.6 ± 0.8 268.9 ± 17.7
(d) (b) + KME Features 0.02 1739.1 1536 1630.2 ± 31.6 55.7 ± 1.1 40.1 ± 0.4 1.9 ± 0.1 348.2 ± 22.7
(e) All Features 0.02 1704.3 1536 1655.9 ± 26.4 63.4 ± 2.8 46.3 ± 3.3 3.8 ± 0.3 363.4 ± 20.6
(f) Final Model 0.005 1825.1 1536 1763.6 ± 21.5 68.4 ± 2.1 53.7 ± 2.2 3.3 ± 0.4 429.8 ± 25.6
IV-B Hardware Design & Sim-to-Real
In this work, we aimed to train a policy in simulation, running on an NVIDIA RTX 4090, and transfer it to real without a policy adaptation phase. During inference, the agent interacts with our real Kinova 6-DOF Jaco 2 through the Kinova ROS API and a custom REST interface, running on a separate low-grade workstation at a 60Hz control rate. Unlike prior works focused on artificial plants with soft leafy resistance, our test branches are randomly chosen from real trees, ensuring sufficient stiffness to resist clearance. A single fixed Intel RealSense D405 camera, operating at 30 fps, provides the RGB-D ($480\times 640$) observation of the workspace, illustrated in Fig. 2(c). The segmented real branch point clouds are constructed from the RGB-D camera images; in contrast, simulation tree points are randomly sampled from the coarse-grained branch segment surfaces. The real robot point cloud is built once from meshes, just as in the simulation, and updated at each step with the 3D coordinate transformation tree from ROS.
Furthermore, to address the significant sim-to-real discrepancies in robot parameters, contact dynamics, and gravity compensation, exacerbated by velocity control, we leverage Segmented Steady State Error Control (SSED)[11]. This scheme alternates between applying an action $a_{t}\sim\pi_{w}(\cdot|s_{t})$ to the desired state $s^{d}_{t}$, instead of the current, but periodically synchronising $s^{d}_{t}$ to $s_{t}$, in effect, rejecting short-term steady-state errors and clearing long-term accumulated offsets.
IV-C Real Vision Pipeline
This section outlines the key transformation steps required to convert the real-time streaming images into RL features. First, we use Grounding DINO [39] to locate the objects in the scene with simple text prompts. The resulting bounding boxes are passed to the Segment Anything SAM-HQ [40] model to generate object masks zero-shot. Segmented masks are then de-projected to the robot base frame, leveraging the depth channel input. A key concern at this stage is that manipulators operating in velocity control mode have a control rate lower bound, 60Hz in our case, required to maintain velocity. Current state-of-the-art grounding and segmentation tools are unable to achieve this without significantly compromising the mask quality. Therefore, we use a video object segmentation framework Cutie [41], to propagate the mask across time-steps, occasionally replacing it with updated masks to prevent quality loss while tracking. Moreover, we use Python Shared Memory modules and asynchronous queues to go beyond the camera-constrained fps and meet the control frequency requirement. Notably, such a pipeline is problem agnostic; specifically, the grounding text prompt alone needs to be modified to switch from the power line context to a cable inspection scenario, a promising direction for future work.
V EXPERIMENTS AND RESULTS
Figure 4: Ablation showing the relevance of key feature groups in our multi-modal policy. (-) indicates the removal of the single specified group. Each data point is a trained policy with a varying noise level. Boxes indicate median and IQR.
Experiment 1: First, we perform a simulation ablation study to justify our feature choices with the setup from Section IV-A. For each combination, we train independent policies by injecting varying levels of Gaussian Noise $\mathcal{N}(0,\sigma^{2})$ into the point clouds with the $3\sigma$ rule, i.e., $\sigma=d_{\text{max}}/3$, $d_{\text{max}}$ being the maximum deviation representative of the camera depth inaccuracy. The ablation results are shown in Fig. 4 and Table II. The box plot data points are policies constructed with $d_{\text{max}}$ levels ranging from $0.005$ to $0.04$, displaying the distribution of test rewards and success rates on a single unseen test set. On the other hand, each row in Table II is an individual policy tested on multiple test sets, listing a variety of additional evaluation metrics, with rows (a)-(e) showing the results applying a median noise $d_{\text{max}}=0.02$, representative of the ablation study. By contrast, row (f) shows our best model with a lower $d_{\text{max}}=0.005$, the depth inaccuracy of the D405 camera used for real experiments.
TABLE III: Representational Efficiency
$\bm{GF_{t}^{(1024)}}$ (secs) $\bm{GF_{t}^{(1)}}$ (secs) Train Time (mins) Train Rew Max Envs
PointNet++ 21.29 0.7706 1487.2 1345.6 1024
Point2Vec 0.204 0.0454 75.2 1379.6 1024
KME 0.0012 0.0012 27.5 1449.9 6144
TABLE IV: Real Results
Branches Attempts Test SR
Branch 1 9 44.4%
Branch 2 26 42.3%
The tabulated evaluation metrics are described as follows: Training Reward (Train Rew) and Test Reward (Test Rew) assess the agent’s performance during training and on unseen test scenarios, respectively. Test Success Rate (Test SR %) quantifies the percentage of test environments (Trials) where all occlusions were removed from the power line. Given that the branch collection will continue to resist clearance even after success is reached, we choose a conservative Test SR, considering a task as a success only if the line is occlusion-free for at least 10 consecutive steps. Occlusion reduction (Occ Drop %) shows the percentage drop in heuristic ($h_{t}$) from the start to end of test episodes averaged across all environments, while the (Touch %) reflects the frequency of undesirable instances where the arm breaches the pre-set safety margin $\delta_{sm}$ around the power line. The final metric (Steps in Succ) measures the efficiency of successful trajectories; as an example, the final row (f) implies that, on average, in a $1000$ step trajectory, the arm starting from the home position, traversed to the occlusion locations and moved the occluding branches away from the line in $1000-429.8=570.2$ steps, while the remaining $429.8$ steps were completely occlusion free for potential pruning, despite continued cluster resistance. All metrics except the power line touch (Touch %) are higher, the better.
The plot in Fig. 4 demonstrates that both the KME feature group capturing the global point cloud attributes and the contact detection classifier form essential constituents of our multi-modal policy, and removal of either would hurt performance. Along the same lines, Table. II shows that despite the stochasticity of individual policies and high noise injection, the progressive addition of feature groups induces improvement across all evaluation metrics. To highlight, we report all test success rates without distinguishing between feasible and infeasible tasks, i.e., all occlusion patterns are randomly generated and may contain scenarios that cannot be solved by a single arm. In some unsolvable instances that contribute to the failures, the arm cannot generate sufficient torques (due to joint limits) to overcome the collective resistance of obstructing branches; in others, multiple disconnected branches can occlude the line at locations not closely spaced, only one of which can be removed at a time.
Experiment 2: Second, we evaluate the representational efficiency of our approach to other point cloud feature extractors, namely, a Pytorch variant [42] of the well-known PointNet++ [35] and the Point2Vec [43]. Specifically, we replace the four global point features from Table I, i.e., $\overline{\mu_{\mathbb{P}}}[P^{wbr}]$, $\overline{\mu_{\mathbb{P}}}[P^{zbr}]$ , $\overline{\mu_{\mathbb{P}}}[P^{clr}]$ and $\overline{\mu_{\mathbb{P}}}[P^{rob}]$ with embeddings built from PointNet++ and Point2Vec. In each case, we use pre-trained checkpoints, down-sample the input points, and employ PCA post-feature extraction wherever necessary for consistency with our approach. We train each set for 250 epochs with 1024 environments, and the corresponding time taken (Train Time) and the accumulated training reward (Train Rew) at the end are tabulated. In addition, we compute ${GF_{t}^{(E)}}$ denoting the average time taken to build the 4x global point features only for a single RL time-step for $E$ environments. $GF_{t}^{(1024)}$ is indicative of the representation efficiency while training on 1024 environments and $GF_{t}^{(1)}$ represents the inference performance for one environment. The results are tabulated in Table IV, where lower is better for the time metrics, such as, $GF_{t}^{(1024)}$, $GF_{t}^{(1)}$, and Train Time while higher values are better for others.
Comparing the time taken for a single time-step and for the overall training, from Table IV, our kernel-based approach shows orders of magnitude improvement, with similar training rewards. For the PointNet++ and Point2Vec cases, the maximum possible environments (Max Envs) we could spin up in our hardware was 1024, an indication of the GPU memory constraint due to the heavy pre-trained checkpoints. Notice that our final KME-based model training, row (f) from Table II, which uses 6x gym environments and 5x iterations compared to this experiment, takes just 4.5 hours to complete on an RTX 4090. Furthermore, the tabulated $GF_{t}^{(1)}$ values indicate that both PointNet++ and Point2Vec cannot meet the 60Hz real-time inference frequency requirement, justifying our approach. However, we acknowledge that improved efficiency may be at the cost of expressivity and that advanced baseline versions or quantisation techniques can reduce this time difference.
Experiment 3: Finally, we go beyond simulation validation to evaluate our approach in a real-world setting under laboratory conditions. In real, we test two branches from different species, modifying either the branch orientation or the power line position in each experiment. Furthermore, in each test, we start with an $h_{t}$ of at least $0.8$, and consider both partial occlusion clearance and any touch to the power line as failure instances. The results are in Table IV.
Corresponding real videos (see supplement and Fig. 3) demonstrate that the simulation-trained policy can adapt zero-shot to arbitrary branch structures with multiple forking patterns despite the presence of point cloud noise and leaf-induced clutter unseen in training. The trajectories exhibit novel characteristics, effectively using the forearm, upper and lower wrists, gripper, and the unactuated end-effector. In some instances, we observe self-reconfiguring strategies to prop up the occluding branches, while in others, the arm inserts itself between the branches and the line to separate them. Further, these results validate our RL-based approach in discovering efficient strategies and adapting to complex scenarios, well beyond what humans can demonstrate.
VI LIMITATIONS AND FUTURE WORK
A few directions for future work to reduce the sim-to-real gap and improve the success rates are as follows. First, the quality of perception can be enhanced with multiple camera views or with partial-to-complete point cloud models for surface reconstruction. Second, to introduce additional real-world complexities during simulation, integrating with visual tree reconstruction works [44] that identifies the stem structure beneath the leaf clutter, or system identification methods [29] that infers parameters governing branch dynamics, may be necessary. Third, external whole-arm tactile sensors [10] could be leveraged instead of our proprioceptive-only model to enhance the real-world applicability and effectiveness of contact-aware policies. Finally, the generalisability of our approach needs to be tested in out-of-context distributions, for example, with cables, ropes, or sutures.